OpenStudy (anonymous):

MEDAL!~ Verify my answer please. What is the area of an isosceles right triangle if the side opposite the right angle is 4 in? my answer: approx 5.66in^2

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

hmm wait a second... I got the wrong 4 =) ok so the hypotenuse is 4

OpenStudy (anonymous):

lol yeah was about to type that <:D

OpenStudy (anonymous):

though, I'm confused. What should I use? Pythagorean or trigo functions? D:>

OpenStudy (anonymous):

and when should I use them? sorry for the questions xD

OpenStudy (jdoe0001):

 \(\bf \cfrac{1}{2}\cdot 2\sqrt{2}\cdot 2\sqrt{2}\implies \cfrac{4\sqrt{4}}{2}\implies 4 \)